OpenStudy (a1234):

Is there anything wrong with this work? cosθ/2 = √((1 + cosθ)/2) cosθ = 24/25 cosθ/2 = √(1 + (24/25)/2 cosθ/2 = √((49/25)/2) cosθ/2 = 7/√50

jimthompson5910 (jim_thompson5910):

Then you can optionally rationalize the denominator by multiplying top and bottom by \(\Large \sqrt{50}\)

jimthompson5910 (jim_thompson5910):

Keep in mind that \[\Large \sqrt{50} = \sqrt{25*2} = \sqrt{25}*\sqrt{2} = 5\sqrt{2}\]

OpenStudy (a1234):

Nothing else is wrong with it?

OpenStudy (a1234):

Because my teacher didn't give me credit for it

jimthompson5910 (jim_thompson5910):

if you rationalize the denominator and use the fact that \[\Large \sqrt{50}=5\sqrt{2}\] then you should get \[\Large \cos\left(\frac{\theta}{2}\right) = \frac{7\sqrt{2}}{10}\]

OpenStudy (a1234):

Ok...thanks!

jimthompson5910 (jim_thompson5910):

no problem